# VAMP-Net: An Interpretable Multi-Path Framework of Genomic Permutation-Invariant Set Attention and Quality-Aware 1D-CNN for MTB Drug Resistance

### v1

@@ 3 Data Description
Data for this study were obtained from Comprehensive Resistance Prediction for Tuberculosis: an International Consortium (CRyPTIC) (cryptic2022data). This globally important resource consists of a robust set of 12,289 clinical isolates of Mycobacterium tuberculosis, collected from 23 countries. Importantly, the dataset provides matched whole-genome sequencing (WGS) information with quantitative Minimum Inhibitory Concentration (MIC) measurements for 13 essential antitubercular drugs, covering the entire treatment landscape: first-line agents (rifampicin, isoniazid, ethambutol), key second-line drugs (amikacin, kanamycin, rifabutin, levofloxacin, moxifloxacin, ethionamide), and newer/repurposed therapeutics (bedaquiline, clofazimine, delamanid, linezolid).
@@
4 Methodology
Our approach introduces a novel Variant-Aware Multi-Path Network (VAMP-Net) that synergistically couples a Set Transformer architecture with a quality-aware 1D-Convolutional Neural Network for genetic variant analysis. This framework is designed as a two-branch fusion model (Figure 1).
The first branch, a Set Transformer, is specifically engineered to process the genetic variant sequence as an unordered set of tokens. This branch employs the Set Attention Block (SAB) to capture the complex non-sequential relationships between the variants, reflecting the inherent permutation-invariant nature of a set.
The second branch of our framework is a quality-aware 1D-CNN. This path focuses on local sequential features. For each variant, it takes as input a corresponding quality score sequence. The 1D-CNN is efficient in capturing the pattern in these local quality signals, thus offering the model critical information on the confidence and reliability of a variant call.
Figure 1: Overview of the Variant-Aware Multi-Path Network (VAMP-Net) Architecture
Finally, the outputs of the Set Transformer and the 1D-CNN branches are fused through a late-fusion mechanism, as pesented in the Figure 1. The representations from both branches are concatenated and passed to a final classification layer. This fusion allows the model to make predictions based on a rich, multi-faceted understanding of the data: the global, set-based relationships between variants from the SAB, and the local, sequential quality signals from the 1D-CNN. This multi-scale learning (local + global) provides both robustness (resistance to noisy data) and biological interpretability (attention highlighting key variants).
To formulate the problem addressed, a genomic sample is represented as a set of $N$ genetic variants: $V=\{v_{1},v_{2},...,v_{N}\}$ where each variant $v_{i}$ is described by $(\textbf{g}_{i},\textbf{f}_{i})$, with $\textbf{g}_{i}$ representing a sequence of genetic alteration and $\textbf{f}_{i}\in\mathbb{R}^{d}$ denoting the associated feature vector extracted from Variant Call Format (VCF) files. The prediction task is binary classification: $y=f_{\theta}(V)$, where $y\in\{0,1\}$ indicates resistance or susceptibility to a given drug, and $f_{\theta}$ is a parameterized deep learning model.
4.1 Data Representation: Genetic Variants as a Set of Tokens
The dataset used in this study comprises genetic variant sequences of Mycobacterium tuberculosis (MTB), as described in Section 3. Each sample is represented by a set of all identified genetic variants, resulting in sequences of variable length.
@@ 4.1 Data Representation: Genetic Variants as a Set of Tokens
Transformer-based architectures have fundamentally transformed the processing of sequential data, exhibiting remarkable ability to capture long-range dependencies and contextual relationships. In the genomics domain, recent transformer models have achieved promising results; however, most existing approaches primarily rely on reference DNA sequences or K-mer tokenization as inputs (consens2025interpreting). While such representations effectively capture local sequence patterns, they often treat variants as abstract tokens embedded within a continuous sequence. This abstraction can obscure the explicit biological relationships and combinatorial interactions among individual genetic alterations—relationships that are essential for understanding complex traits such as drug resistance.
@@
To address this limitation, we propose a variant-centric representation that explicitly encodes both the molecular identity and genomic position of each variant as a single, biologically meaningful token. In this framework, each token directly represents a specific reference-to-alternate allele substitution, enabling the model to learn variant-level interactions rather than merely sequence-level dependencies. This transition from sequence-level attention to variant-level interaction attention provides a more interpretable and biologically grounded foundation for modeling genotype–phenotype associations. Specifically, we employ an attention mechanism that operates on a ChromPos_Ref>AltChromPos\_Ref>Alt representation, designed to capture biologically meaningful variant–variant interactions. In this representation, each genetic variant $\textbf{a}_{k}$ is uniquely identified by its chromosome, precise genomic position, reference allele, and alternate allele (e.g., chr1:12345:A>>G). This compact and canonical identifier serves as a semantically rich token, allowing each sample’s genomic profile $\textbf{g}_{i}$ to be expressed as a discrete set of variant tokens T: $\textbf{g}_{i}=\{a_{1},a_{2},...,a_{T}\}$. Such a representation enables the transformer to reason directly over meaningful biological entities—facilitating the discovery of high-order, non-linear relationships among variants that underlie phenotypic outcomes such as drug resistance.
This representation is particularly well-suited for a permutation-invariant modeling framework. By adopting this symbolic encoding, the entire collection of genetic variants for each sample can be treated as a set rather than an ordered sequence. This design allows the model to learn complex, non-linear dependencies and interactions among variants independent of their order in the dataset or genome file. The proposed tokenization strategy therefore serves as a key enabler of our set-based learning paradigm, marking a distinct shift from conventional sequence-oriented approaches and focus on the intrinsic relationships between variants themselves.
@@ 4.2 Model Architecture: Variant-Aware Multi-Path Network (VAMP-Net)
Variant-Aware Multi-Path Network (VAMP-Net), a dual-path deep learning framework specifically designed to address the limitations of conventional single-stream models in genomic data analysis. VAMP-Net is architected to process two distinct sources of information essential for robust variant-based prediction ( Figure 1): (1) the symbolic identity and biological context of each genetic variant and (2) the quality metrics of quantitative sequencing that reflect data reliability.
@@
The framework is structured as follows:
@@ 4.2.1 Path-1: Set Attention Transformer Block (SAB) on Genetic Variants
Figure 2: Path-1 Architecture: Set Attention Transformer Block (SAB) for Symbolic Variant Processing
@@
The first contribution adapts the Set Transformer framework (lee2019set) of the unmasked SAB for modeling genetic variant sequences represented as unordered tokens. Unlike standard transformers for sequential data, no attention masks —neither causal masks is applied, ensuring each variant can attend to every other variant without positional or causal constraints, nor padding masks is applied. This model serves as a theoretically pure baseline for evaluating the efficacy of a purely permutation-invariant approach. The input to the model is a batch of padded sequences, where each variant is represented by its token embedding. Within the SAB, the self-attention mechanism computes attention scores for every token pair, including those involving the meaningless padding tokens. The model learns to encode the relationships between all tokens, with no explicit instruction to ignore the padded ones. This approach tests the model’s capacity to learn a meaningful representation of the data despite the presence of noisy, non-informative tokens. The training objective is to perform a specific task of Drug Resistance, based on their set of variants. The hyperparameters for this model, including the number of SAB layers, hidden dimensions, and learning rate, are provided in the Experimental Section 5.4 .
@@ 4.2.2 Path-2 : Quality-Aware 1D-CNN
Figure 3: Path-2 Architecture: Quality-Aware 1D Convolutional Neural Network (1D-CNN) for VCF Feature Processing
@@ 4.2.3 Multi-Path Fusion Mechanism
The final stage of the VAMP-Net architecture involves the integration of the learned representations from the two specialized paths. The outputs of the two branches—the SAB’s permutation-invariant set representation ($\mathbf{z}_{\text{SAB}}$) from Path-1, and the 1D-CNN’s quality-aware feature vector ($\mathbf{z}_{\text{CNN}}$) from Path-2 —are concatenated to form a comprehensive, fused representation $\mathbf{z}_{\text{fused}}$ (Equation 8):
$\mathbf{z}_{\text{fused}}=[\mathbf{z}_{\text{SAB}};\mathbf{z}_{\text{CNN}}]$ (8)
This late-fusion approach is crucial, as it allows each branch to specialize independently in its unique data modality (symbolic variant topology vs. quantitative feature sequence) before their insights are combined. The fused vector is then passed through a fully connected network to a final output layer for the prediction of drug resistance.
To determine the optimal strategy for combining these disparate signals, we investigate three distinct modulation mechanisms between the output of Path-1 (transformer) and Path-2 (1D-CNN):
@@
    Suppression (De-emphasis of Quality):
    The first fusion mechanism is inspired by LSTM’s forget gate control. The purpose of this gate in LSTM is to control how much of the old state is remembered in the new state. By taking the sigmoid of the current input $C_{t}$ and the hidden state $H_{t-1}$, values are transformed into the probability domain, and then multiplied element-wise with $C_{t-1}$, the past internal state.
    We modify the purpose and formula of this gate to suit our case: by taking the sigmoid of Path-2, we use the additional features to control the confidence of the current variants, and then multiply the result by the output of Path-1. The Suppression fusion is defined in Equation 9, where $\odot$ denotes element-wise multiplication.
    $\mathbf{g}=\sigma(\mathbf{z}_{\text{CNN}})$
@@
    This formulation applies a multiplicative suppression to the output of Path-1 ($\mathbf{z}_{\text{SAB}}$). Since the modulation score $\mathbf{g}$ is constrained to $[0,1]$, the resulting fused vector $\mathbf{z}_{\text{fused}}$ will have element values ranging between 0 and the corresponding values in $\mathbf{z}_{\text{SAB}}$. In effect, Path-2’s quality features provide a scalar factor that selectively scales the influence of each variant, and if the quality metrics indicate low confidence, the variant’s representation can be driven toward 0, effectively performing a dynamic, data-driven form of variant feature selection during the fusion process.
@@
    Amplification (Emphasis on Quality):
    The second investigated fusion strategy, Amplification, is designed to explicitly test the hypothesis that quality and low-level features (Path-2) provide a critical signal gain to the core symbolic variant representation (Path-1). This mechanism ensures that high-quality, high-confidence variants can have proportionally increase in impact, thereby emphasizing a robust subset of the data.
    Unlike the Suppression mechanism, this formulation guarantees that the symbolic representation $\mathbf{z}_{\text{SAB}}$ is always passed through to the final output, regardless of the quality metrics. The Quality-Aware vector $\mathbf{z}_{\text{CNN}}$ acts as an additive gain control. We adapt the gating mechanism as defined in Equation 10, where $\odot$ denotes element-wise multiplication.
    $\mathbf{g}=\sigma(\mathbf{z}_{\text{CNN}})$
@@
    This formulation achieves a soft integration that preserves the full contribution of the core variant signal. By introducing the constant $1$ into the scalar factor $(1+\mathbf{g})$, the mechanism guarantees a baseline preservation, ensuring that the fused output $\mathbf{z}_{\text{fused}}$ is never less than the original symbolic representation $\mathbf{z}_{\text{SAB}}$. As the modulation score $\mathbf{g}$ is constrained to the range $[0,1]$, the total scalar factor $(1+\mathbf{g})$ remains bounded within the interval $[1,2]$. This results in a controlled form of additive gain, where the most robustly supported variants—those with high-quality metrics—receive a maximum scaling of $200\%$ of their original magnitude. This approach avoids the risk of zeroing out potentially relevant variants, allowing all identified mutations to participate in the final prediction while prioritizing those with the highest associated confidence.
@@
    Adaptive Amplification & Suppression
    The final fusion strategy is Adaptive Amplification & Suppression, designed to provide the model with a dynamic, context-aware control over the Path-1 signal. This mechanism combines the feature selection capability of Suppression with the signal enhancement capability of Amplification, allowing the influence of the symbolic variant representation ($\mathbf{z}_{\text{SAB}}$) to be either scaled down or scaled up based on the quality metrics ($\mathbf{z}_{\text{CNN}}$).
    We synthesize the previous two modulation strategies by modifying the gating factor. Using the same modulation score $\mathbf{g}=\sigma(\mathbf{z}_{\text{CNN}})$, the fused vector is calculated as Equation 11.
    $\mathbf{z}_{\text{fused}}=(2\odot\mathbf{g}-1)\odot\mathbf{z}_{\text{SAB}}$ (11)
    where $\odot$ denotes element-wise multiplication.
    This adaptive mechanism creates a bipolar scaling factor which introduces a powerful, new operational dynamic. As the modulation score $\mathbf{g}$, derived from the quality features, is constrained to the range $[0,1]$, the resulting scaling factor $(2\odot\mathbf{g}-1)$ is dynamically mapped to the closed interval [1,1][-1,1]. This allows the model to continuously transition between Suppression Zone (noise filtering) and Amplification Zone (signal gain) based on the input quality. Specifically, when quality features are poor ($\mathbf{g}\approx 0$), the scaling factor approaches 1-1, resulting in a negative scaling of the variant representation, effectively acting as an aggressive noise filter by reversing the polarity of low-confidence signal. Conversely, when quality features are excellent ($\mathbf{g}\approx 1$), the factor approaches $1$, leading to maximum positive scaling and thus prioritizing the most robustly supported variants. This ability to span both positive and negative scalar values allows the model to not only ignore noise but to actively penalize uncertain variants, providing a comprehensive, context-aware control signal to the final prediction layers .

### v2

@@ 3.3.1 Data Description
Data for this study were obtained from Comprehensive Resistance Prediction for Tuberculosis: an International Consortium (CRyPTIC) [9]. This globally important resource consists of a robust set of 12,289 clinical isolates of Mycobacterium tuberculosis, collected from 23 countries. Importantly, the dataset provides matched whole-genome sequencing (WGS) information with quantitative Minimum Inhibitory Concentration (MIC) measurements for 13 essential antitubercular drugs, covering the entire treatment landscape: first-line agents (rifampicin, isoniazid, ethambutol), key second-line drugs (amikacin, kanamycin, rifabutin, levofloxacin, moxifloxacin, ethionamide), and newer/repurposed therapeutics (bedaquiline, clofazimine, delamanid, linezolid).
@@
3.3.2 Data Representation: Genetic Variants as a Set of Tokens
@@
Transformer-based architectures have fundamentally transformed the processing of sequential data, exhibiting remarkable ability to capture long-range dependencies and contextual relationships. In the genomics domain, recent transformer models have achieved promising results; however, most existing approaches primarily rely on reference DNA sequences or K-mer tokenization as inputs [8]. While such representations effectively capture local sequence patterns, they often treat variants as abstract tokens embedded within a continuous sequence. This abstraction can obscure the explicit biological relationships and combinatorial interactions among individual genetic alterations, relationships that are essential for understanding complex traits such as drug resistance.
@@
To address this limitation, we propose a variant-centric representation that explicitly encodes both the molecular identity and genomic position of each variant as a single, biologically meaningful token. In this framework, each token directly represents a specific reference-to-alternate allele substitution, enabling the model to learn variant-level interactions rather than merely sequence-level dependencies. This transition from sequence-level attention to variant-level interaction attention provides a more interpretable and biologically grounded foundation for modeling genotype–phenotype associations. Specifically, we employ an attention mechanism that operates on a ChromPos_Ref>AltChromPos\_Ref>Alt representation, designed to capture biologically meaningful variant–variant interactions. In this representation, each genetic variant $\textbf{a}_{k}$ is uniquely identified by its chromosome, precise genomic position, reference allele, and alternate allele (e.g., chr1:12345:A>>G). This compact and canonical identifier serves as a semantically rich token, allowing each sample’s genomic profile $\textbf{g}_{i}$ to be expressed as a discrete set of variant tokens T: $\textbf{g}_{i}=\{a_{1},a_{2},...,a_{T}\}$. Such a representation enables the transformer to reason directly over meaningful biological entities, facilitating the discovery of high-order, non-linear relationships among variants that underlie phenotypic outcomes such as drug resistance.
To mitigate out-of-vocabulary (OOV) effects, we employ a data-driven tokenization scheme in which a BERT WordPiece tokenizer is trained on the complete set of unique variant strings extracted from the dataset As illustrated in Figure 2. The tokenizer operates without lowercasing and learns a compact vocabulary composed of genomic primitives and frequently occurring subpatterns, including numeric positions, canonical nucleotides (A, C, G, T), and structural delimiters (e.g., >> and ::). At inference time, each variant string is decomposed into valid subword units using a greedy longest-match strategy, ensuring that even rare or previously unseen variant compositions within the same representational schema can be expressed through combinations of known tokens. This approach substantially reduces reliance on the $[UNK]$ token and preserves the compositional structure of genomic variants. The resulting sub-token embeddings are subsequently aggregated to form a unified high-dimensional representation of each variant, which is then provided as input to the downstream Path-1 Set Attention Transformer. By aligning the tokenizer with the generative structure of the data, this method enables robust modeling of the long-tail distribution of genomic variants while maintaining vocabulary efficiency.
Figure 2: WordPiece-based tokenization strategy for genomic variant encoding. This figure details the mechanism used to mitigate Out-of-Vocabulary (OOV) errors during the processing of non-canonical mutations. An input variant (e.g., $761155:C>T$) is processed via WordPiece-based tokenization, which decomposes the symbolic string into fundamental sub-units: numerical subunits, nucleotides, and delimiters. By leveraging Fixed Vocabulary Tables that contain these universal genomic components, VAMP-Net can construct unique embedding vectors for previously unseen rare coordinates. This approach avoids the information loss associated with traditional $[UNK]$ (unknown) tokens. The resulting Sub-token Embeddings are aggregated to produce a single, high-dimensional Variant Embedding Vector ($V_{761155:C>T}$), which is then passed to the Path-1 Set Attention Transformer. This method ensures that the model can learn from the "long tail" of genomic data while maintaining a compact and efficient vocabulary.
@@ 3.4 Model Architecture: Variant-Aware Multi-Path Network (VAMP-Net)
Variant-Aware Multi-Path Network (VAMP-Net), a dual-path deep learning network specifically designed to address the limitations of conventional single-stream models in genomic data analysis. VAMP-Net is architected to process two distinct sources of information essential for robust variant-based prediction ( Figure 1): (1) the symbolic identity and biological context of each genetic variant and (2) the quality metrics of quantitative sequencing that reflect data reliability.
@@
The model is structured as follows:
@@ 3.4.1 Path-1: Set Attention Transformer Block (SAB) on Genetic Variants
Figure 3: Path-1 Architecture: Set Attention Transformer Block (SAB) for Symbolic Variant Processing. This pathway facilitates the extraction of high-order epistatic features from raw genomic data. Raw variant text inputs (formatted as Position_Ref>AltPosition\_Ref>Alt) are first mapped into a continuous vector space via a Compositional Embedding layer. These embeddings are processed through a series of NN stacked Set Attention Blocks (SAB). Unlike standard Transformers, the SAB utilizes self-attention mechanisms without positional encodings, ensuring the model remains permutation-invariant. Within each SAB, Multi-Head Attention (MHA) computes Query (Q), Key (K), and Value (V) interactions to model the non-linear dependencies between distant mutations. The final output is a set of per-variant feature representations that encode the biological topology of the isolate, specifically capturing the synergistic effects of multiple mutations on the drug resistance phenotype.
@@
The first contribution adapts the Set Transformer framework [25] of the unmasked SAB for modeling genetic variant sequences represented as unordered tokens. Unlike standard transformers for sequential data, no attention masks —neither causal masks is applied, ensuring each variant can attend to every other variant without positional or causal constraints, nor padding masks is applied. This model serves as a theoretically pure baseline for evaluating the efficacy of a purely permutation-invariant approach. The input to the model is a batch of padded sequences, where each variant is represented by its token embedding. Within the SAB, the self-attention mechanism computes attention scores for every token pair, including those involving the meaningless padding tokens. The model learns to encode the relationships between all tokens, with no explicit instruction to ignore the padded ones. This approach tests the model’s capacity to learn a meaningful representation of the data despite the presence of noisy, non-informative tokens. The training objective is to perform a specific task of Drug Resistance, based on their set of variants. The hyperparameters for this model, including the number of SAB layers, hidden dimensions, and learning rate, are provided in the Experimental Section 4.2.3.
@@ 3.4.2 Path-2 : Quality-Aware 1D-CNN
Figure 4: Path-2 Architecture: Quality-Aware 1D Convolutional Neural Network (1D-CNN) for VCF Feature Processing. This pathway is dedicated to the quantitative assessment of sequencing reliability. The input consists of a multi-channel tensor of VCF-derived features, including Genotype Calls (GT), Read Depth (DP), and Fraction of Supporting Reads (FRS), among others listed in the Data Callout Board. These features are processed through a series of NN stacked Convolutional Blocks, each containing a 1D-convolutional layer for local pattern extraction. Unlike traditional binary filters, this 1D-CNN learns a continuous, high-dimensional representation of technical confidence for each variant. These per-variant feature representations are then utilized by the fusion module to adaptively weight the biological signals from Path-1, ensuring that resistance predictions are grounded in high-quality sequencing data.
@@ 3.4.3 Multi-Path Fusion Mechanism
The final stage of the VAMP-Net architecture involves the formal integration of the learned representations from the two specialized paths: the biological topology manifold and the technical quality manifold.
- Formal Integration Framework:
Let $\mathbf{z}_{\text{SAB}}\in\mathbb{R}^{d}$ be the permutation-invariant set representation from Path-1, and $\mathbf{z}_{\text{CNN}}\in\mathbb{R}^{k}$ be the quality-aware feature vector from Path-2. To ensure that technical confidence can modulate biological signal, we define a Gated Fusion Module that maps technical noise patterns to a modulation intensity vector $\mathbf{g}\in\mathbb{R}^{d}$ via a sigmoid-activated transformation:
$\mathbf{g}=\sigma(\mathbf{W}g.\mathbf{z}_{\text{CNN}}+\mathbf{b}_{g})$ (8)
where $\mathbf{W}_{g}$ and $\mathbf{b}_{g}$ are learnable parameters. This gating vector acts as a "Technical Auditor," determining the trustworthiness of each dimension in the biological feature space.
- Fusion Modulation Strategies:
To determine the optimal strategy for combining these disparate signals, we investigate three distinct modulation mechanisms that resolve conflicts between biological data and technical noise:
@@
    Suppression (De-emphasis of Noise):
    Inspired by the "forget gate" in Long Short-Term Memory (LSTM) networks, this mechanism performs element-wise multiplication to scale the biological signal:
@@
    As $\mathbf{g}\in[0,1]$, this formulation performs a dynamic, data-driven feature selection. If Path-2 identifies poor sequencing quality (e.g., low depth), $\mathbf{g}\to 0$, effectively suppressing the influence of that variant to prevent the propagation of sequencing artifacts.
@@
    Amplification (Emphasis on Signal): To test if high-quality data provides a critical signal gain, we introduce a baseline-preserving mechanism:
@@
    This ensures that the symbolic representation $\mathbf{z}_{\text{SAB}}$ is always preserved, while high-confidence variants receive a gain of up to $200\%$. This prioritizes robustly supported mutations without risking the "zeroing out" of potentially relevant signals.
@@
    Adaptive Amplification and Suppression (Proposed Best):
    Our final strategy synthesizes both approaches into a bipolar scaling factor that provides the model with dynamic, context-aware control:
    $\mathbf{z}_{\text{fused}}=(2\mathbf{g}-1)\odot\mathbf{z}_{\text{SAB}}$ (11)
    By mapping the scaling factor to the interval [1,1][-1,1], the model can continuously transition between a Suppression Zone (noise filtering) and an Amplification Zone (signal gain). This allows the model to not only ignore noise but to actively penalize uncertain variants by reversing their signal polarity, mirroring the rigorous audit-trail logic used by clinical microbiologists.
Final Classification
The resulting vector $\mathbf{z}_{\text{fused}}$ is passed through a global average pooling layer and a Softmax-activated dense layer to produce the final resistance probability $P(y=1)$:
$P(y=1)=\text{Softmax}(\mathbf{W}f.\mathbf{z}_{\text{fused}}+\mathbf{b}_{f})$ (12)
- Functional Interpretation of Adaptive Fusion:
The bipolar scaling mechanism in VAMP-Net is designed to resolve conflicts between biological signals and technical noise. This mathematical operation corresponds to two distinct operational modes:
  • Adaptive Suppression: If Path-2 detects poor sequencing quality (e.g., low depth or high genotype noise), the gating factor $\mathbf{g}$ approaches 0, and the scaling factor becomes negative. This effectively filters the variant out, preventing the model from assigning resistance based on a potential sequencing artifact.
  • Adaptive Amplification: Conversely, when Path-2 confirms high technical confidence for a non-canonical mutation, the scaling factor approaches $1$. This amplifies the signal, allowing the model to prioritize novel, high-confidence epistatic interactions that might be missed by static filters.
This design ensures that biological causality is always contextualized by data reliability, mirroring the audit-trail logic used by clinical microbiologists during manual isolate review.